OpenStudy (anonymous):

Give an example of a rational function that has no horizontal asymptote and a vertical asymptote at x = 1.

OpenStudy (anonymous):

@amistre64 @phi @mathstudent55

OpenStudy (amistre64):

what are some basic rules that you can apply here?

OpenStudy (anonymous):

Im really not sure. I was going to approach this with guess and check but I figured you could help me approach it

OpenStudy (amistre64):

there are 3 fundamental rules for HA calculations when the degree of the top is _________ the degree of the bottom, the HA is ______

OpenStudy (amistre64):

degrees can only be greater than, equal to, or less than .... each case gives us a specific HA assessment

OpenStudy (anonymous):

Okay. Im not positive as to what you mean

OpenStudy (amistre64):

your material for your course should define 3 cases for rational function

OpenStudy (amistre64):

by comparing the degree of the top and bottom polynomials, the horizontal asymptote can be determined

OpenStudy (anonymous):

Okay. So is the a formula of some type that shows me the relation of the two?

OpenStudy (amistre64):

recall that a HA is only defined for very large values of x therefore the leading terms of the top and bottom polys take control of the values\[\frac{ax^n}{bx^m}\]

OpenStudy (amistre64):

if n=m, then for large values of x, the asymptote is a/b when n<m then the bottom gets bigger quicker than the top and we go to zero when n>m then the top get bigger quicker and we ahve no HA

OpenStudy (amistre64):

any of this not make sense?

OpenStudy (anonymous):

I think I follow

OpenStudy (amistre64):

compare: \[\frac{2x^3}{3x^3}\] x^3/x^3=1 so, HA = 2/3 \[\frac{2x}{3x^3}\] 10000/10000^3 so we have a very small value that approaches zero \[\frac{2x^3}{3x}\] 10000^3/10000 this is just a bigger and bigger number

OpenStudy (amistre64):

we know what the bottom needs to be right?

OpenStudy (anonymous):

bottom = 2/3?

OpenStudy (amistre64):

what defines a vertical asymptote?

OpenStudy (anonymous):

The distance of the curve from the x axis

OpenStudy (amistre64):

a vertical asymptote is a factor that cannot be removed .... \[\frac{something}{x-1}\] as long as something doesnt have x-1 as a factor, then the bottom cant be removed and we have a VA at x=1

OpenStudy (amistre64):

now, in order for the thing to have no HA, the top has to be a poly of greater degree than the bottom. any ideas for the top?

OpenStudy (anonymous):

x^2?

OpenStudy (amistre64):

x^2 is fine :)

OpenStudy (anonymous):

Okay Great :)

OpenStudy (anonymous):

So what next?

OpenStudy (amistre64):

submit the rational expression of course

OpenStudy (anonymous):

So its just x^2/x-2

OpenStudy (anonymous):

I mean x-1

OpenStudy (amistre64):

x-1 but yes

OpenStudy (anonymous):

Oh wow! That was simpler than I thought. Thanks so much!!

OpenStudy (amistre64):

yw